A 0.7 Liter aqueous solution oh HCIO4 has a density of 1.67 g/mL. What is the molarity of the HCIO4 solution?
Answer #1

Solution:

Density of HClO4 = 1.67 g/ml = 1670 g/L

Volume = 0.7 L

Then, mass (m) of HClO4 = Volume x Density

m = 0.7 L x 1670 g L -1 = 1169 g

Since, molarity is defined as,

Molarity = mass / (molar mass x Volume in L)

Molar mass of HClO4 =100.46 g/mol, therefore,

Molarity = 1169 g / (100.46 g mol-1 x 0.7 L) = 15.12 M

Therefore, molarity = 15.12 M